A certain bright star has a temperature of 10,600 K, and is believed to have a radius of order 70 times the radius of the Sun. At what orbital radius would a planet have a flux of radiation comparable to that on our Earth – ie – habitable like the Earth? At this stellar temperature, what colour does the star have?
Answer #1

a) For this problem, we can consider the star as a perfect black body.

So, the energy emitted by it per unit area is

If the planet has a radius r, the total power emitted by the planet is

If the planet is at a distance R from the star, the flux received by the planet per unit area is

We know that T = 10,600 K

r = 70*rsun = 70*696*10^6 = 4.872*10^10 m

For earth, F = 1388 W/m^2

So,

b) The wavelength of radiation is given by

This light will be in the UV region.

So, the color of the star as it appears to human eye will be blue.